OpenStudy (anonymous):

Calculus: I have a test in 4 hours and I need to learn this stuff quick. If you could lay out the steps and a breif examplantion so I can follow your rationale that would be great. 3. Find the following indefi nite integrals. If you use integration by substitution, write what you chose to be u and find du. Don't forget to add the constant! (a) {integral sign} [x-root(3x2 + 1)] dx (b) {integral sign} [(4+lnx)^2/x] dx (c) {integral sign} [(x^2e^(6-x^3)] dx (d) {integral sign} [(2x^3-3x)/(x^4-3x^2+1)] dx

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

for number 1 is that \[\int x\sqrt{3x^2 + 1}dx\]

OpenStudy (lgbasallote):

@MercerC you here?

OpenStudy (anonymous):

Yes it is

OpenStudy (lgbasallote):

okay...here's a hint..try letting \[u = 3x^2 + 1\] what do you get as du then?

OpenStudy (anonymous):

nevermind I think I'm just going to repeat this class. this is overwhelming

OpenStudy (lgbasallote):

i was just doing integration by substitution

OpenStudy (lgbasallote):

to find du you just take the derivative of \(3x^2 + 1\)

OpenStudy (anonymous):

6x

OpenStudy (lgbasallote):

right. \[du = 6xdx\] so \[xdx = \frac{du}{6}\] agree? i just divided both sides by 6

OpenStudy (anonymous):

okay

OpenStudy (lgbasallote):

so now..our integral was \[\int x\sqrt{3x^2 + 1} dx\] but i can also rearrange it like \[\int (\sqrt{3x^2 + 1})xdx\] right? commutative property of multiplication?

OpenStudy (anonymous):

okay

OpenStudy (lgbasallote):

so...remember i said \[u = 3x^2 + 1\] and \[\frac{du}{6} = xdx\] so i can change the integral to \[\int (\sqrt{3x^2 + 1} )xdx \implies \int (\sqrt u) \frac{du}{6}\] following so far?

OpenStudy (anonymous):

yep

OpenStudy (lgbasallote):

now..i'll take 1/6 out of the integral \[\frac 16 \int (\sqrt u) du\] then change \(\sqrt u\) into \(u^{1/2}\) \[\implies \frac 16 \int (u^{1/2})du\] okay?

OpenStudy (anonymous):

following...

OpenStudy (lgbasallote):

good...now do you remember the rule in integral that looks like this.. \[\Large \int x^n \text dx \implies \frac{x^{n+1}}{n+1}\] does that seem familiar?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

I'm sorry, I'm going to stop you. Thank you for your help. But I think there is just way too much of this test that I don't know.

OpenStudy (lgbasallote):

this is called the power formula..it is used a lot when integrating...basically when you have an integral that looks like \[\int x^n dx\] the integral of that is just \[\frac{x^{n+1}}{n+1}\] x is the variable and n is the exponent. in this problem we have \[\int u^{1/2} du\] so the integral of that is \[\large \frac{u^{\frac 12 + 1}}{\frac 12 + 1}\] make sense?

OpenStudy (anonymous):

an accellerated calculus class was a bad idea

OpenStudy (lgbasallote):

do you have background knowledge in calculus?

OpenStudy (anonymous):

nope, its the only course I need to take to graduate.why is my solution this:

OpenStudy (anonymous):

\[1/9(3x ^{2}+1)^{3/2}+C\]

OpenStudy (lgbasallote):

well like i was saying \[\Large \frac{u^{\frac 12 + 1}}{\frac 12 + 1} \implies \frac{u^{\frac 32}}{\frac 32} \implies \frac{2u^{\frac 32}}{3}\] then you had a 1/6 outside so \[\Large \frac 16 \times \frac{2u^{\frac 32}}{3} \implies \frac 13 \times \frac{u^{\frac 32}}{3} \implies \frac{u^{\frac 32}}{9}\] then you substitute u back into \(u = 3x^2 + 1\) \[\Large \frac{u^{\frac 32}}{9} \implies \frac{(3x^2 + 1)^{\frac 32}}{9}\] sorry if it looks complicated :(

OpenStudy (anonymous):

It's okay. I'm sure its easy for you. I'm just having a tough time understanding what the heck I'm learning. It's all so abstract and not relevant to any of my education. I wish I would have taken this 4 years ago when suggested.

OpenStudy (lgbasallote):

believe it or not...it took me two tries to pass the course...that's why i know it well now..

OpenStudy (anonymous):

Me too, this accelerated course didn't really teach me much. It was just an overflow of information and terminology that I don't understand.

OpenStudy (anonymous):

right now, I'll probably get a 15% on the final. So, I'm thinking of calling it a night and trying next semester.

OpenStudy (anonymous):

thanks for your help.

OpenStudy (lgbasallote):

actually...i passed my second try in calculus because of this site :D this site taught me a lot in calculus back in the days